# Using Enriched Category Theory to Construct the Nearest Neighbour Classification Algorithm

### v1

@@ 3 Constructing The Nearest Neighbours Algorithm
This section explores the construction of the Nearest Neighbours Algorithm, given a dataset of points in a metric space, and classification labels, using Enriched Category Theory. Starting with a dataset of $n$ pairs $(x_{0},y_{0}),...,(x_{n},y_{n})$, the $x_{i}$ values are elements of a metric space $X$, and the $y_{i}$ values are class labels. Given a new point $x\in X$, what is the correct class label to associate with it?
@@
From the format of the dataset, the primary characteristic of the data points are the distances between them. This would suggest that the natural choice for the enriched categories are Lawvere metric spaces, i.e. Cost enriched categories. The first step is to find an appropriate representation of the data. An individual data point, $(x_{i},y_{i})$, has three components. An index value $i$, an associated point in the metric space $x_{i}$, and the classification label $y_{i}$. The $n$ index values can be stored in a Cost-enriched category $N$. The metric space $X$ can clearly also be represented as a Cost-enriched category $X$, but the class labels can also be represented in a similar way, as the contents of the Cost-enriched category $Y$, which contains all of the possible class labels. With these categories, the information of the dataset can be represented by two functors. $F:N\rightarrow X$ maps the index values to their associated position in the metric space $x_{i}$. The functor $T:N\rightarrow Y$, similarly, maps data indexes to class labels.
Though it is now clear what objects the various enriched categories contain, it remains to determine what the hom-objects of each category should be. In the case of the metric space $X$, it is clear that between any points $a,b\in X$, the hom object $X(a,b)$ should correspond directly with the distance metric on X. It is less clear what the choice should be for the categories $N$ and $Y$.
@@
Proceeding with the intuition that the hom-objects, or in this case the distances, between objects should encode meaningful information about the data, the objects of $N$, the indexes, possess no explicit relation to each other. This would suggest that the distances between indexes should be as ”un-constraining as possible”. In the context of enriched categories, the lack of constraint would suggest that the Functors from $N$ to any other Cost category, should correspond directly with maps from the objects of $N$ to the other category. To achieve this, the category $N$ can be given the discrete metric, shown in the following equation.
$N(i,j)=\begin{cases}0&i=j\\ \infty&i\neq j\end{cases}$
@@
Recalling that functors between Cost-Categories are distance non-increasing functions, the discrete metric means that this condition is trivially satisfied, as the objects of $N$ are as distant from each other as possible. This models the lack of a relationship between the data indexes. The same logic can be applied to the objects of $Y$. Class labels should also have no meaningful relation to each other, so the discrete metric can be applied to $Y$ as well. With the categories $N$, $X$, $Y$ and the functors $F$, and $T$, the dataset can be represented by the following diagram.
@@
To find the classes of all the points in $X$ would optimistically be to find a suitable candidate for the dotted arrow from $X$ to $Y$. However, there is an issue. It is expected that two classification regions in $X$ may be touching, producing a boundary between classification regions which can have a trivially small distance. If we insist that classes are assigned by functors, then the functors must be distance non increasing. This would require that the classes in $Y$ have a distance of zero from each other. It is tempting to think that one should not assign $Y$ the discrete metric, but this has an unfortunate consequence. Within the language of Enriched Category Theory, the hom-objects are the only way to distinguish between objects of a category. Setting all of the distances between objects in $Y$ to zero would make all of the classes indistinguishable from each other in any categorical construction. It was correct to assign $Y$ the discrete metric, but not to expect the classifications to be represented by a functor. The classifications can in fact be represented by a profunctor $NNA:X\nrightarrow Y$.
With the expectation that the correct classification is represented by a profunctor, we can attempt to produce this profunctor directly by composition. The functors $F$ and $T$ both have two canonical profunctors associated with them. By selecting these profunctors appropriately, we can compose them to produce a profunctor from $X$ to $Y$. This can be done with the profunctors $F^{*}:X\nrightarrow N$ and $T_{*}:N\nrightarrow Y$.
@@
As previously discussed, the profunctor $F^{*}:X\nrightarrow N$ measures the distance between a point in $X$ and the image of a data point in $N$. The profunctor $T_{*}:N\nrightarrow Y$ does something similar, but because it is produced by a functor between discrete categories, its outputs are even easier to interpret. If a data index $i$ has a class $y$, i.e. $Ti=y$, then $T_{*}(y,i)$ will be 00. However, if $i$ does not have class $y$ then $T_{*}(y,i)$ is infinity. Substituting these profunctors into the profunctor composition formula produces the following equation.
$(T_{*}\circ F^{*})(y,x)=\inf_{i\in N}\ (F^{*}(i,x)+T_{*}(y,i))$
The interpretation of this composition is relatively straight forward. If the class of $i$ selected by the infimum is not $y$, then $T(y,i)$ will be infinity, making the entire sum as large or larger than any other possible value. However, if the $i$ selected was of class $y$, then the formula returns $\inf_{i\in N}F^{*}(i,x)$. In other words the composition $(T_{*}\circ F^{*})(y,x)$ returns the distance from $x$ to the closest data point which is of class $y$. This could also be interpretted as evaluating the infimum of a partition of the indexes which have the class y  [^4] Note that the following expression re-uses the notation $NT(y)$ introduced in section 2.1 to represent the partition subset of $N$ with classes $y$, $NT(y)=\{i\in N\ |\ Ti=y\}$.
$(T_{*}\circ F^{*})(y,x)=\inf_{i\in NT(y)}d(Fi,x)$
A useful outcome, but not quite the NNA. There is one additional step. In order to reproduce the NNA we need to compare the output of the profunctor $T_{*}\circ F^{*}$, to a similar composition with a profunctor that has no knowledge of the classes, $\mathbf{1}_{NY}:N\nrightarrow Y$.
$(\mathbf{1}_{NY}\circ F^{*})(y,x)=\inf_{i\in N}\ (F^{*}(i,x)+\mathbf{1}_{NY}(y,i))=\inf_{i\in N}F^{*}(i,x)$
Given a point $x\in X$ and class $y\in Y$, the profunctor $(\mathbf{1}_{NY}\circ F^{*})(y,x)$ gives the distance to the closest point in the dataset (i.e. in the image of $F$). This composition has forgotten all class information. Finally, to reconstruct the NNA classification it only remains to compare the outputs of both profunctors. As their outputs are objects of the Cost category, the natural comparison is their hom-object in Cost.
$NNA(y,x):=Cost((\mathbf{1}_{NY}\circ F^{*})(y,x),\ (T_{*}\circ F^{*})(y,x)):X\nrightarrow Y$
Because the arrows in Cost encode the ordering information, this leads to the expression:
$NNA(y,x)=(\mathbf{1}_{NY}\circ F^{*})(y,x)\ \geq\ (T_{*}\circ F^{*})(y,x)$
A point $x$ is taken to have class $y$ when $NNA(y,x)$ is true. Consider the situation that the closest data point $Fj$ to $x$ has class $y$, then $T(y,j)=0$. The left hand side of the inequality finds the smallest distance from x to a data point with any class and the right hand side finds the smallest distance to a data point with class y. When the closest data point to $x$ has class $y$, the left hand side returns the same value as the right hand side and the inequality is true.
$NNA(y,x)$ $\quad\Leftrightarrow\quad Cost(\quad(\mathbf{1}_{NY}\circ F^{*})(y,x)\ ,\quad(T_{*}\circ F^{*})(y,x)\quad)$
$\quad\Leftrightarrow\quad(\mathbf{1}_{NY}\circ F^{*})(y,x)\quad\geq\quad(T_{*}\circ F^{*})(y,x)$
$\quad\Leftrightarrow\quad\inf_{i\in N}F^{*}(i,x)\quad\geq\quad\inf_{i\in N}\ (F^{*}(i,x)+T(y,i))$
$\quad\Leftrightarrow\quad F^{*}(j,x)\quad\geq\quad F^{*}(j,x)+T(y,j)$
$\quad\Leftrightarrow\quad F^{*}(j,x)\ \geq\ F^{*}(j,x)$
$\quad\Leftrightarrow\quad True$
$NNA(y,x)$ $\quad\Leftrightarrow\quad Cost(\quad(\mathbf{1}_{NY}\circ F^{*})(y,x)\ ,\quad(T_{*}\circ F^{*})(y,x)\quad)\quad$
$\quad\Leftrightarrow\quad(\mathbf{1}_{NY}\circ F^{*})(y,x)\quad\geq\quad(T_{*}\circ F^{*})(y,x)$
$\quad\Leftrightarrow\quad\inf_{i\in N}F^{*}(i,x)\quad\geq\quad\inf_{i\in N}\ (F^{*}(i,x)+T(y,i))$
$\quad\Leftrightarrow\quad\inf_{i\in N}F^{*}(i,x)\quad\geq\quad\inf_{i\in NT(y)}F^{*}(i,x)$
$\quad\Leftrightarrow\quad\inf_{i\in N}F^{*}(i,x)\quad=\quad\inf_{i\in NT(y)}F^{*}(i,x)$
$\quad\Leftrightarrow\quad\inf_{i\in N}X(Fi,x)\quad=\quad\inf_{i\in NT(y)}X(Fi,x)$
$\quad\Leftrightarrow\quad\inf_{i\in N}d(Fi,x)\quad=\quad\inf_{i\in NT(y)}d(Fi,x)$
4 Future Work
Given the diversity of Machine Learning algorithms, and the natural generalising power of Enriched Category Theory, there are numerous avenues to explore for future extensions of this work.
Firstly, the construction of the NNA in section 3 does not require any specific properties of Cost-enriched categories to define. This leads very naturally to a candidate definition of the V-enriched Nearest Neighbours Algorithm (V-NNA).
@@ 4 Future Work
This immediately begs the question of whether this definition has useful properties in other bases of enrichment. Though the previous section interpretted the hom-object of the base of enrichment in its non enriched form for the sake of clarity, future works would benefit from considering the self enriched form of the hom-object. In the case of the Cost-NNA, interpreting the hom-object with truncated subtraction rather than an inequality.
@@
Researchers who aren’t interested in Machine Learning would possibly consider the Voronoi diagram as a more interesting outcome of the V-NNA. By assigning each index its own separate class, the NNA partitions the metric space dependent on each individual point rather than each individual class. In this instance, the partitions generated in other bases of enrichment may prove interesting.
It is also interesting to ask what other algorithms can be presented in this language. An obvious generalisation of the Nearest Neighbours Algorithm is the k Nearest Neighbours Algorithm, where classification of a point is based on the majority classification of the k closest points to it. Beyond this, there are many ML algorithms which depend purely on the distance metric of their dataset, so many may also be found as constructions of Cost-enriched categories.

### v2

@@ 3 Constructing The Nearest Neighbours Algorithm
Starting with a dataset of $n$ pairs $(x_{1},y_{1}),...,(x_{n},y_{n})$, the $x_{i}$ values are elements of a metric space $X$, and the $y_{i}$ values are class labels. Given a new point $x\in X$, what is the correct class label to associate with it?
@@
An individual data point, $(x_{i},y_{i})$, has three components. An index value $i$, an associated point in the metric space $x_{i}$, and the classification label $y_{i}$. The $n$ index values, the points, and the class labels can be stored in the Cost-enriched category $N$, $X$, and $Y$. The information of the dataset can be represented by two functors. $F:N\rightarrow X$ maps the index values to their associated position in the metric space $x_{i}$. The functor $T:N\rightarrow Y$ maps data indexes to class labels.
@@
In the case of the metric space $X$, it is clear that between any points $a,b\in X$, the hom object $X(a,b)$ should correspond directly with the distance metric on $X$. The objects of $N$, the indexes, possess no explicit relation to each other. This would suggest that the distances between indexes should be as un-constraining as possible. The lack of constraint would suggest that the Functors from $N$ to any other Cost category, should correspond directly with maps from the objects of $N$ to the other category. To achieve this, the category $N$ can be given the discrete metric.
$N(i,j)=\begin{cases}0&i=j\\ \infty&i\neq j\end{cases}$ (12)
@@
As the objects of $N$ are maximally distance, any mapping of objects from $N$ satisfies the condition of being distance non increasing. The same logic can be applied to the objects of $Y$. Class labels should also have no meaningful relation to each other, so the discrete metric can be applied to $Y$ as well. The complete dataset can be represented by the following diagram.
${X}$${N}$${Y}$$\scriptstyle{T}$$\scriptstyle{F}$ (13)
To find the classes of all the points in $X$ would be to find a suitable candidate for the dotted arrow from $X$ to $Y$. Such an extension problem models the regression of $T$ on $X$. Finding a relation from dependent variables to independent variables. However, it is expected that two classification regions in $X$ may be touching, producing a boundary between classification regions which can have a trivially small distance. If the classes are assigned by functors, then the functors must be distance non-increasing. This would require that the classes in $Y$ have a distance of zero from each other, which would have an unfortunate consequence. The hom-objects are the only way to distinguish between objects of a category. Setting all of the distances between objects in $Y$ to zero would make all of the classes indistinguishable from each other in any categorical construction. It was correct to assign $Y$ the discrete metric, but not to expect the classifications to be represented by a functor. The classifications can in fact be represented by a profunctor $\textit{NNA}:X\nrightarrow Y$.
The functors $F$ and $T$ both have two canonical profunctors associated with them. By selecting these profunctors appropriately, they may be composed to produce a profunctor from $X$ to $Y$.
@@
The profunctor $F^{*}:X\nrightarrow N$ measures the distance between a point in $X$ and the image of a data point in $N$. The profunctor $T_{*}:N\nrightarrow Y$ does something similar, but because it is produced by a functor between discrete categories, its outputs are even easier to interpret. If a data index $i$ has a class $y$, i.e. $Ti=y$, then $T_{*}(y,i)$ will be 00. However, if $i$ does not have class $y$ then $T_{*}(y,i)$ is infinity. Substituting these profunctors into the profunctor composition formula produces the following equation.
$(T_{*}\circ F^{*})(y,x)=\inf_{i\in N}\ (F^{*}(i,x)+T_{*}(y,i))$ (15)
If the class of $i$ selected by the infimum is not $y$, then $T(y,i)$ will be infinity, making the entire sum as large or larger than any other possible value. However, if the $i$ selected was of class $y$, then the formula returns $\inf_{i\in N}F^{*}(i,x)$. In other words, the composition $(T_{*}\circ F^{*})(y,x)$ returns the distance from $x$ to the closest data point which is of class $y$. This could also be interpreted as evaluating the infimum of a partition of the indexes which have the class $y$  [^4] Note that the following expression re-uses the notation $NT(y)$ introduced in section 2.1 to represent the partition subset of $N$ with classes $y$, $NT(y)=\{i\in N\ |\ Ti=y\}$.
$(T_{*}\circ F^{*})(y,x)=\inf_{i\in NT(y)}d(Fi,x)$ (16)
In order to reproduce the NNA the profunctor $T_{*}\circ F^{*}$ needs to be compared to a similar composition with a profunctor that has no knowledge of the classes, $\mathbf{1}_{NY}:N\nrightarrow Y$.
$(\mathbf{1}_{NY}\circ F^{*})(y,x)$ $=\inf_{i\in N}\ (F^{*}(i,x)+\mathbf{1}_{NY}(y,i))$ (17)
$=\inf_{i\in N}F^{*}(i,x)$ (18)
Given a point $x\in X$ and class $y\in Y$, the profunctor $(\mathbf{1}_{NY}\circ F^{*})(y,x)$ gives the distance to the closest point in the dataset (i.e. in the image of $F$). This composition has forgotten all class information. Comparing the outputs of both profunctors produces the NNA. As their outputs are objects of the Cost category, the natural comparison is their hom-object in Cost.
NNA $:X\nrightarrow Y$ (19)
NNA $(y,x):=Cost((\mathbf{1}_{NY}\circ F^{*})(y,x),\ (T_{*}\circ F^{*})(y,x))$ (20)
Because Cost can be viewed as a Bool enriched category, and therefore a preorder, this leads to the expression:
$\textit{NNA}(y,x)=(\mathbf{1}_{NY}\circ F^{*})(y,x)\ \geq\ (T_{*}\circ F^{*})(y,x)$ (21)
A point $x$ is taken to have class $y$ when $\textit{NNA}(y,x)$ is true. Consider the situation that the closest data point $Fj$ to $x$ has class $y$, then $T(y,j)=0$. The left hand side of the inequality finds the smallest distance from x to a data point with any class and the right hand side finds the smallest distance to a data point with class y. When the closest data point to $x$ has class $y$, the left hand side returns the same value as the right hand side and the inequality is true.
$\textit{NNA}(y,x)$ $\Leftrightarrow Cost(\ (\mathbf{1}_{NY}\circ F^{*})(y,x),\ (T_{*}\circ F^{*})(y,x)\ )$ (22)
$\Leftrightarrow(\mathbf{1}_{NY}\circ F^{*})(y,x)\quad\geq\quad(T_{*}\circ F^{*})(y,x)$ (23)
$\Leftrightarrow\inf_{i\in N}F^{*}(i,x)\quad\geq\quad\inf_{i\in N}\ (F^{*}(i,x)+T(y,i))$ (24)
$\Leftrightarrow F^{*}(j,x)\quad\geq\quad F^{*}(j,x)+T(y,j)$ (25)
$\Leftrightarrow F^{*}(j,x)\ \geq\ F^{*}(j,x)$ (26)
$\Leftrightarrow True$ (27)
$\textit{NNA}(y,x)$ $\Leftrightarrow Cost(\ (\mathbf{1}_{NY}\circ F^{*})(y,x)\ ,\ (T_{*}\circ F^{*})(y,x)\ )$ (28)
$\Leftrightarrow(\mathbf{1}_{NY}\circ F^{*})(y,x)\quad\geq\quad(T_{*}\circ F^{*})(y,x)$ (29)
$\Leftrightarrow\inf_{i\in N}F^{*}(i,x)\quad\geq\quad\inf_{i\in N}\ (F^{*}(i,x)+T_{*}(y,i))$ (30)
$\Leftrightarrow\inf_{i\in N}F^{*}(i,x)\quad\geq\quad\inf_{i\in NT(y)}F^{*}(i,x)$ (31)
$\Leftrightarrow\inf_{i\in N}F^{*}(i,x)\quad=\quad\inf_{i\in NT(y)}F^{*}(i,x)$ (32)
$\Leftrightarrow\inf_{i\in N}X(Fi,x)\quad=\quad\inf_{i\in NT(y)}X(Fi,x)$ (33)
$\Leftrightarrow\inf_{i\in N}d(Fi,x)\quad=\quad\inf_{i\in NT(y)}d(Fi,x)$ (34)
4 Generalising the Nearest Neighbours Algorithm
4.1 K Nearest Neighbours
The $k$-NNA is an extension of the NNA that bases its output on an aggregate classification formed from the classes of the $k$ nearest neighbours. To produce this generalisation, the construction must store the information associated with $k$ neighbouring data points, and a choice needs to be made concerning the method of aggregating the classes of these points. Given a Cost enriched category $C$, one can produce the $k$ times tensor product of $C$ with itself, using the tensor product introduced in section 2.3.
$C^{k}:=C_{1}\otimes...\otimes C_{k}\quad\quad C=C_{i}$ (35)
The objects of $C^{k}$ are $k$-tuples of the objects of $C$. The tensor product operates on the hom objects of $C$ by applying the monoidal product of the base of enrichment. In the case of Cost, the monoidal product is addition.
$C^{k}(\vec{x},\vec{y})=C_{1}(\vec{x}_{1},\vec{y}_{1})\otimes...\otimes C_{k}(\vec{x}_{k},\vec{y}_{k})$ (36)
The tensor product of functors maps tuples to tuples, with each component functor acting element-wise on the tuples.
$F^{k}(\vec{x})=(F(\vec{x}_{1}),...,F(\vec{x}_{k}))$ (37)
Because the tensor product of enriched categories acts on both categories and functors, one can produce the $k$ times tensor product of the diagram which defines the dataset.
@@ 4.1 K Nearest Neighbours
This diagram can store the information required for the $k$-NNA classification. However, it includes tuples that would usually be excluded, i.e., where a particular data point appears multiple times. Usually, the $k$ nearest neighbours are $k$ distinct neighbours. By including tuples which duplicate data points, the algorithm can base its aggregate classification on a tuple of duplicate points. To avoid this issue, we can introduce a subcategory of $N^{k}$, which only includes tuples with no duplicate points. This subcategory is equipped with a functor which is an inclusion on objects, mapping the subcategory into the original space, $I:\overline{N^{k}}\rightarrow N^{k}$. The introduction of $I$ allows the diagram to be restricted, by composing $I$ with $F$ and $T$, such that the algorithm only acknowledges tuples of distinct neighbours.
@@
The second step in this procedure is to correct the outputs and inputs of the diagram by composing profunctors that convert $k$-tuples back into individual points of $X$ and $Y$. The $k$ times tensor product of $X$ with itself comes equipped with a profunctor functor that identifies the $k$-tuples formed from a single element, $\Delta:X\nrightarrow X^{k}$ The profunctor identifies such tuples by comparing a single element of $X$ with each element of the tuple using the distance metric of $X$.
$\Delta(\vec{x},x)=X(\vec{x}_{1},x)\otimes...\otimes X(\vec{x}_{k},x)$ (40)
When every element of $\vec{x}$ is zero distance from $x$ then $\Delta(\vec{x},x)=0$. There are multiple viable options for the profunctor used to correct the outputs of the diagram. The aggregator profunctor takes the form $A:Y^{k}\nrightarrow Y$. Part of the variability in its definition is how $A$ handles ties, where the most common labels are equally prevalent amongst the $k$ neighbours. Whenever there is a clash, a greedy aggregation policy may assign both labels, a conservative policy would assign no labels, and a biased policy may always prefer one label over another. Ultimately, $A$ becomes a hyperparameter. The commonality between these schemes is that $A$ is invariant to permutations in the order of the elements, that it is 00 when a tuple of neighbours does possess the given classification label, and it is $\infty$ otherwise. The following diagram may be produced with access to the profunctors $\Delta$ and $A$.
@@
In parallel with the strategy of section 2.1 this allows the $k$-NNA algorithm to be defined through profunctor composition.
$k$-NNA $:X\nrightarrow Y$ (42)
$k$-NNA $:=A\circ Cost((\mathbf{1}_{\overline{N^{k}}Y^{k}}\circ(FI)^{*}),\ ((TI)_{*}\circ(FI)^{*}))\circ\Delta$ (43)
The definition of $k$-NNA can also be presented more concisely with reference to the original NNA algorithm, constructed over the functors $F^{k}I$ and $T^{k}I$. For clarity, we notate the original NNA construction as $\textit{NNA}_{F,T}$.
[…]
Presenting the $k$-NNA using the NNA construction will also allow the arguments of section 3 to be leveraged without having to repeat them. Each component profunctor must yield true for the composite $k$-NNA to yield true. In the case of Cost-enriched profunctors, the output 00 is interpretted interchangeably with the value true. To convince ourselves that the $k$-NNA is behaving as expected, we may follow a point $x\in X$ around the diagram.
[…]
In summary, $k$-NNA can be produced using the NNA algorithm by duplicating the original dataset with a $k$ fold tensor product. A choice of class label aggregation policy allows $k$-tuples of labels to be reduced to a single class label.
4.2 Label Metrics for Soft Boundaries and Dependent Classifications
To demonstrate that ECT can extend the utility of an existing algorithms it can be shown that the construction of the NNA begins exhibiting novel and potentially valuable behaviours when its construction parameters are allowed to vary outside of what is usually possible.
The particular variation explored in this section occurs through the definition of the space of classification labels, $Y$. In the traditional implementation, the collection of classification labels is a set of values with no additional structure; as the algorithm is usually implemented directly from its description, it is unclear how one would incorporate additional structure into its decision-making. Section 3 shows that when $Y$ is considered a discrete category, the ECT construction of NNA reproduces precisely the expected behaviour. However, $Y$ is not required to be a discrete category in general. By framing the construction within the category of Lawvere metric spaces, the enriched NNA can accept $Y$ as any Lawvere metric space. Incorporating information about the similarity of classification labels as a metric allows the enriched classifier to produce softer classification boundaries. This could be formed from a semantic similarity metric for word-based labels or a more traditional metric for vector-based labels. Metrics associated with $Y$, the category of classification labels, will be referred to as label metrics.
Label metrics are incorporated into the classification decision through $T_{*}$ as $T_{*}(y,i)=Y(y,Ti)$ When the label metric is discrete, $T_{*}$ serves only to identify when points are identical. However, for label metrics with bounded distances, the enriched NNA may present a point as having a particular classification if that label is within a certain distance of the label of the nearest neighbour. Consider, for example, a function from the unit interval to the unit interval. A sample of points from this function can be presented with functors $F$ and $T$, allowing the application of the NNA. Both $X$ and $Y$ values can be given the standard metric of the real numbers rather than forcing $Y$ to be discrete. The outputs of the NNA and $4$-NNA for every $X$ and $Y$ position are visualised in Fig 2.
Figure 2: A plot of the values of NNA (left) and $4$-NNA (right) for $x,y\in[0,1]$. $30$ points were uniformly sampled from the interval $[0,1]$ and transformed by the function $f(x)=0.4+0.1\sin(10x)-0.7x^{2}+0.7x^{3}$ then randomly scaled by ±$\pm 5\%$ to produce the $Y$ values. The hom objects where taken to be $X(a,b)=Y(a,b)=|b-a|$. The aggregation policy chosen for $4$-NNA is only zero when all tuple components agree on the class. Both colour scales clip outputs outside their stated range.
The images demonstrate the output values of the profunctors, showing how they evaluate the nearness to truth of each coordinate point. Operating in this fashion, the NNA begins to act more as a regression model than as a classifier.
The presentation of $Y$ as a Lawvere metric space also allows it to adopt an asymmetric metric. This induces what one might refer to as dependent classifications, where the assignment of one label may depend on the assignment of another. For example, consider the classification labels Dog, Cat, Bird, and Mammal. The label metric between Dog, Cat, and Bird can be taken to be discrete. As all Dogs and Cats are Mammals, the relationship can be represented by allowing the metric to become zero when evaluated in one direction. $Y(\textit{Mammal},\textit{Dog})=Y(\textit{Mammal},\textit{Cat})=0$. As not all Mammals are Dogs or Cats, the converse relationship is assigned an infinite distance. $Y(\textit{Dog},\textit{Mammal})=Y(\textit{Cat},\textit{Mammal})=\infty$. As birds are not mammals and mammals are not birds, then these labels are an infinite distance from each other. $Y(\textit{Bird},\textit{Mammal})=Y(\textit{Mammal},\textit{Bird})=\infty$. Assume for some value $x$, that $\textit{NNA}(x,\textit{Dog})=True$, then the two component profunctors of the NNA are equal. $(\mathbf{1}_{NY}\circ F^{*})(\textit{Dog},x)=(T_{*}\circ F^{*})(\textit{Dog},x)$. The statement that $Y(\textit{Mammal},\textit{Dog})=0$ allows it to be inserted into the definition of the right-hand side to demonstrate that $(\mathbf{1}_{NY}\circ F^{*})(\textit{Dog},x)=(T_{*}\circ F^{*})(\textit{Mammal},x)$.
$(\mathbf{1}_{NY}\circ F^{*})(\textit{Dog},x)$ $=(T_{*}\circ F^{*})(\textit{Dog},x)$ (45)
$=\inf_{i\in N}\ (F^{*}(i,x)+T_{*}(\textit{Dog},i))$ (46)
$=\inf_{i\in N}\ (F^{*}(i,x)+Y(\textit{Dog},Ti))$ (47)
$\begin{split}&=\inf_{i\in N}\ (F^{*}(i,x)+Y(\textit{Mammal},\textit{Dog})\\ &\quad\quad+Y(\textit{Dog},Ti))\end{split}$ (48)
$\geq\inf_{i\in N}\ (F^{*}(i,x)+Y(\textit{Mammal},Ti))$ (49)
$\geq(T_{*}\circ F^{*})(\textit{Mammal},x)$ (50)
The presence of the infimum in the definition of $(\mathbf{1}_{NY}\circ F^{*})(\textit{Dog},x)$ forces the inequality to be an equality. By the definition of the left-hand side, we know that $(\mathbf{1}_{NY}\circ F^{*})(\textit{Dog},x)=(\mathbf{1}_{NY}\circ F^{*})(\textit{Mammal},x)$ as $(\mathbf{1}_{NY}\circ F^{*})$ is invariant to the first input. Together, these statements induce the following equality.
$(\mathbf{1}_{NY}\circ F^{*})(\textit{Mammal},x)=(T_{*}\circ F^{*})(\textit{Mammal},x)$ (51)
This infers that $\textit{NNA}(\textit{Mammal},x)=True$. Proving the aforementioned dependent classification.
$\textit{NNA}(\textit{Dog},x)\implies\textit{NNA}(\textit{Mammal},x)$ (52)
The same logic also induces a dependent classification between Cat and Mammal. However, as the distance from Mammal to Dog or Cat is infinite, neither depends on the Mammal label. Similarly, as the labels Mammal and Bird are mutually infinite distance apart, neither depends on the other.
Exploring the enriched NNA as $Y$ is allowed to take non-discrete and asymmetric metrics demonstrates that the resultant behaviours may apply to a broad range of use cases. When $Y$ is allowed to be non-discrete, the NNA can make regression and classification decisions reminiscent of kernel density algorithms. When $Y$ is allowed to be asymmetric, it can encode dependent relationships reminiscent of simple ontologies.